how would i convert one eighth and one ninth into a decimal number rounded to the nearest thousandth?
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

1/8
just do long division 8 into 1.0000
1/9
9 into 1.00000

Answer by Alan3354(69443)

how would i convert one eighth and one ninth into a decimal number rounded to the nearest thousandth?
--------------
1/8 = 0.125
1/9 = 0.11111....
-------------Add
= 0.236
-----------
Or
1/8 + 1/9 = 9/72 + 8/72 = 17/72
=~ 0.236
The first way is easier, you can use a calculator